An insurance company crashed four cars in succession at 5 miles per hour. The cost of repair for each of the four crashes was $420, $465, $407, $222 . Compute the range, sample variance, and sample standard deviation cost of repair. Compute the mean, median, and mode cost of repair. Compute the mean, median, and mode cost of repair. The mean cost of repair is $__ The median cost of repair is $__ The mode cost of repair is $__ (Round to the nearest cent as needed.)
Answer #1

Mean cost of repair = (420 + 465 + 407 + 222)/4 = $ 378.50

Median cost of repair = Average of 2nd and 3rd term = $ (222 + 407)/2 = $ 314.5

There is no mode.